OpenStudy (lilmike234):

What are the equations of the asymptotes for this hyperbola? x^2/25-y^2/4=1

OpenStudy (zzr0ck3r):

The asymptotes of a hyperbola \(\dfrac{(x - h)^2}{a^2} - \dfrac{(y - k)^2}{b^2} = 1\) are given by \(y - k = \pm \dfrac{b}{a}(x - h)\)

OpenStudy (zzr0ck3r):

For you \(h=0, k=0, a=2, b=5\)

OpenStudy (lilmike234):

SSo what would be my equations?